# Implicit State Estimation via Video Replanning

### v1

@@ 4.2 Rejection-based Replanning
At each planning round, instead of generating a single rollout, the Video Plan Generator produces $N$ candidate plans $\{v^{p}_{1},v^{p}_{2},\dots,v^{p}_{N}\}$ by conditioning on $N$ independently sampled state embeddings from the Retrieval Module. Because each sampled embedding represents a different hypothesis of the environment parameters, the resulting video plans can differ substantially, covering a range of possible outcomes.
@@
where $K$ denotes the distance function. Empirically, we found that using an $L2$ distance directly on the pixel space of video frames works well. The rejection module then selects the plan with the largest distance to past failures,
@@ 4.3 Action Module
The final step is to convert a selected video plan into executable low-level actions for the robot. We follow the idea of dense object- or wrist-tracking from AVDC Ko et al. (2024): the predicted video frames are used to track either the object of interest or, in tasks such as bar picking, the robot’s wrist when object tracking is insufficient. The resulting trajectory is then translated into control commands through simple heuristic policies such as grasp strategies or trajectory-following controllers. This training-free approach enables direct execution of video plans without requiring an additional learned policy.
@@ 5 Experiment
5.1 Meta-World system identification benchmark
Existing manipulation benchmarks primarily emphasize single-episode task completion while assuming full observability, overlooking the rich information available through interaction. Consequently, they fail to capture challenges posed by environmental variation, such as determining whether a door should be pushed or pulled, that humans frequently encounter in real world and can solve with ease. Therefore, in this work, we propose the Meta-World System Identification Benchmark to evaluate online adaptation under unknown system parameters $\theta$. It combines an offline experience dataset with environments exhibiting hidden parameters that agents must infer through interaction. Table 1 shows the set of 5 tasks, including object manipulation (e.g., push bar, pick bar) and puzzle solving (e.g., turn faucet), test adaptability to randomized centers of mass, uncertain interaction modes, and external forces.
@@ 5.1 Meta-World system identification benchmark
Table 1: Meta-World System Identification Benchmark. Each task involves interaction-based inference of system parameters $\theta$, which are either continuous or discrete with varying mode counts.
Task Identification Type
Push Bar Continuous
Pick Bar Continuous
Slide Brick Continuous
Open Box Discrete, 2 modes
Turn Faucet Discrete, 2 modes
@@
Turn faucet. In this task, the robot must rotate a faucet. The faucet’s direction of rotation is either clockwise or counterclockwise.
In the above five tasks, the system parameters cannot be inferred without physical interaction. The details of the tasks are provided in Appendix B.
@@ 5.2 Evaluation metrics
Second, instead of only evaluating task success, we also assess the effectiveness of the replanning mechanism by comparing the replanned videos to groundtruth interaction videos. To quantify this similarity, we use standard video comparison metrics: PSNR, SSIM, and LPIPS Zhang et al. (2018).
Figure 4: Qualitative results of adapted video plans. We show qualitative examples of adaptation trials for each task in the Meta-World Identification Benchmark to illustrate the effectiveness of the proposed method. The framework generates new plans based on previously failed interaction and video plans.
5.3 Implementation and baselines
We evaluate our framework on the 5 tasks in the Meta-World System Identification Benchmark. For our model, the Video Plan Generator is trained on the videos in the offline experience dataset as described in 4.1. Each task is evaluated over 400 trials. We compare our approach with the original AVDC implementation for video planning which does not online adapt, variants of our method that use different context encoders $E$, and other reinforcement learning baselines.
  • AVDC (Ko et al., 2024). The baseline with only the Action Module, i.e., without the Retrieval Module and Rejection Module introduced in this paper.
  • Ours. Our proposed method with $E$ based on DINOv2 feature as described in 4.1.
  • Ours (Refine). A variant of our method with refining state embeddings from scratch as mentioned in 4.1.
  • BC uses a state embedding-conditioned behavior cloning framework where the Policy takes in state and state embedding directly. The state embedding is obtained with $E$ based on DINOv2.
  • CQL (Kumar et al., 2020a) is a state embedding-conditioned actor-critic framework. The Q-function takes in state, action, and state embedding, while the Policy takes in state and state embedding, which is obtained with $E$ based on DINOv2.
@@ 5.4 Results
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
AVDC 19.426 0.700 0.271
Ours (R3M) 19.632 0.704 0.261
Ours (CLIP) 19.584 0.703 0.262
Ours (DINOv2) 19.817 0.708 0.255
Our variants. Among the variations of our method, using the Rejection Module (Ours) alone to generate state embeddings achieves the best overall performance. In contrast, the refinement process introduced in Section 4.1 alone, Ours (Refine), yields similar or slightly worse performance but does not assume data availability at test time, i.e., do not require access to the dataset used to train the diffusion model. Additionally, its formulation as a learned model offers the potential for improved generalization. Finally, we show qualitative results for adaptations in Figure 4.
Video plan evaluation. In Table 3, we report the evaluations of video plans generated by the baseline (AVDC) and our methods with various context encoders $E$, including R3M Nair et al. (2022), CLIP Radford et al. (2021), and DINOv2 Oquab et al. (2024). Our method produces significantly higher-quality videos compared to AVDC. Our method with DINOv2 outperforms its counterpart using R3M or CLIP, validating our design choices.
Figure 5: Rejection and retrieval modules. Both the Rejection and Retrieval Modules improve the performance from AVDC (Aciton Module only) and combining them (ours) yields the best performance. In this figure, the numbers of replans are normalized aggregated among the five tasks.
Figure 6: Rejection module distance. We compare using raw-pixel distance and DINOv2 embedding distance to reject and the former yields better performance. The numbers of replans are normalized and aggregated among the five tasks.
Figure 7: Number of generated video plans. We experiment with generating different numbers of video plans $n\in[1,5]$ and set $n=2$. In this figure, the numbers of replans are normalized and aggregated among the five tasks.
@@ 5.5 Ablation studies
To investigate the effectiveness of the Rejection and Retrieval Modules, we compare AVDC (Action Module only), AVDC+Rejection, AVDC+Retrieval, and our method (with both Rejection and Retrieval). Notably, the AVDC+Rejection baseline is conceptually similar to a recent work Li et al. (2025), which also uses a generator to sample candidate plans and a separate module to select the most viable one given previous interaction attempts. The result in Figure 7 shows the normalized numbers of replans across all the tasks and individual task performance is shown in Appendix D. Both AVDC+Rejection and AVDC+Retrieval outperform AVDC, justifying the effectiveness of the Rejection and Retrieval modules. Our full method (with both the Rejection and Retrieval modules) achieves the best performance, validating the combinations.
@@
We compare pixel-wise distance and DINOv2-based similarity for selecting video plan candidates, as described in Section 4.2. Figure 7 shows that pixel-wise distance performs better, likely because DINOv2 struggles to distinguish these plans due to the nature of its pretraining data.
@@
To study how the number of video plans generated to be rejected affects the performance, we vary the number from 1 to 5 and show the result in Figure 7. The result aggregates the performance across all the tasks. From $n=1$ (AVDC+Retrieval) to $n=2$ (ours), around $10\%$ of replans are saved. However, from $n=2$ to $n=5$, the numbers of replans until success are near the same. An increase in video plan candidates generated by the video generator results in an increase in the computational cost. Therefore, since the performances of $n\in\{2,3,4,5\}$ are comparable, we set $n=2$.
Rejection module distance metric.
To study the impact of the distance function used in the Rejection Module, we evaluated several alternatives. The results show that raw-pixel distance consistently achieves superior or comparable performance relative to the DINOv2 baseline. Based on this finding, we adopt raw-pixel distance as the default choice of $K$ for the Rejection Module.
Table 4: Rejection module distance metric. We compare the performance of using different distance functions for the Rejection Module.
Metric Push Bar Pick Bar Slide Brick Open Box Turn Faucet All (Normalized)
[…]
5.6 Evaluation on real world data
@@
Figure 8: Real-world setup. A WidowX arm attempts to open a door that can either be pulled or pushed open.
We collected 20 real-world videos (10 successes, 10 failures) to train the video model. We address the challenge of training on such a small dataset by adopting a joint training with simulation data: 40% of samples came from real-world data and 60% from the Meta-World System Identification Benchmark. To reduce overfitting to the first frames, which often leads to monotonic plans, we injected Gaussian noise into the first-frames during inference. Furthermore, we set $n=4$ to promote diverse plan generation.
For simplicity, we evaluate the single-episode replanning success rate, conditioned on an initial failure—different from the main evaluation metric. A perfect random policy, which assumes knowledge of the correct modes (e.g., push or pull) but does not use past interaction, would succeed with 50% probability. Our method achieved 15/20 successes, compared to 8/20 for a naive video planner without using past information, demonstrating the effectiveness of our pipeline.
Figure 9: Real-world qualitative results. Given the first frame (a), the robot first attempts to push the door (b). Then, it gets stuck and fails to open the door (b). Finally, it generates a plan that pushes the door instead (c) by leveraging information from previous failure.
Figure 9 shows the qualitative rollout for our real-world experiment. We followed most of the hyperparameters as used in the Meta-World System Identification Benchmark. Additionally, we changed $n=4$ and injected Gaussian noise of standard deviation 0.35 into first-frames with pixel values normalized to $[0,1]$ to mitigate the first-frame bias issue.
@@ 6 Conclusion
We presented a video planning framework that adapts to interaction-time uncertainties by implicitly encoding past experiences into the planning process. Our method filters failed plans, updates models online, and achieves robust adaptation without explicit system identification. Experiments on the Meta-World System Identification Benchmark confirm improved robustness over video-based and RL baselines, highlighting the potential of video as a foundation for adaptive decision-making in uncertain environments.
@@
This work was supported by the Industrial Technology Research Institute, Taiwan, and the National Science and Technology Council, Taiwan (114-2628-E-002 -021-). Shao-Hua Sun was supported by the Yushan Fellow Program by the Ministry of Education, Taiwan.
Reproducibility Statement
We are committed to reproducibility and will release the full codebase for reproducing the experiments introduced in this work. Our experiments were conducted on standard GPUs (V100/3090) with moderate training times, and we include full details in the appendix to facilitate independent verification.
@@ A.1 Website and Code
We present extensive qualitative results and release our code in our website.
@@ Appendix B Benchmark Details of Simulator Tasks
The environments of the five tasks are shown in Figure 10.
B.1 Push Bar
The objective of this task is to push a bar towards the target. Physically, the bar is combined with a mass point and a mass-free bar. Therefore, if the robot does not push the bar exact at the center of mass, torque is imposed on the bar, resulting in the deflection of the bar.
The bar has a length of 0.4 meters. In the pretraining stage the video generation model, the distances between the center of mass and geometric center of the bar are shown in Table 5.
Table 5: Center of mass positions. The distances between the center of mass and the geometric center of the bar in the pretraining stage of the video generation model. The sign represents the direction.
-0.18 -0.165 -0.15 -0.135 -0.12
-0.105 -0.09 -0.075 -0.06 -0.05
-0.03 -0.015 0.0 0.015 0.03
0.045 0.06 0.075 0.09 0.105
0.12 0.135 0.15 0.18
The visualization of this task is shown in Figure 16.
B.2 Pick Bar
The objective of this task is to grasp a bar towards the target. Physically, like push-bar mentioned above, the bar is combined with a mass point and a mass-free bar. Therefore, if the robot does not grasp the bar exact at the center of the mass, torque is imposed on the bar, resulting in the tilt of the bar.
The properties of the bar in this task are the same as the bar in push-bar. And the distances between the center of mass and geometric center of the bar in the pretraining stage are also the same as above.
The visualization of this task is shown in Figure 16.
B.3 Slide Brick
Note that the wooden slope has a relatively low friction coefficient and that the red slope has a relatively high friction coefficient. Therefore, the brick can slide down and accelerate in the wooden slope and decelerate in the red slope. The friction coefficient of the wooden slope is fixed and is set to 0, while the friction coefficient of the red slope is various, and those in the pretraining stage are shown in Table 6.
The visualization of this task is shown in Figure 16.
Table 6: Friction coefficients. The friction coefficients of the red slope in the pretraining stage of the video generation model.
0.24 0.25 0.26 0.27 0.28 0.3 0.32
0.34 0.35 0.36 0.38 0.39 0.4
B.4 Open Box
The visualization of the task is shown in Figure 16.
B.5 Turn Faucet
The visualization of the task is shown in Figure 16.
@@ D.1 State Embedding Feature Space
Figure 17 visualizes the feature space introduced in Section 4.1, using DINOv2 (Oquab et al., 2024) as the encoder. The plots are generated by applying t-SNE to sets of state embeddings from each environment, where for continuous modes we randomly sample four values for visualization. Our method yields a well-structured feature space in which different modes can be separated without additional data-processing, facilitating more effective downstream learning.
@@
D.2 Training Data and Scaling Effects.

### v2

@@ 4.2 Rejection-based Replanning
At each planning round, instead of generating a single rollout, the Video Plan Generator produces $N$ candidate plans $\{v^{p}_{1},v^{p}_{2},\dots,v^{p}_{N}\}$ by conditioning on $N$ independently sampled embeddings from the Retrieval Module. Because each sampled embedding represents a different hypothesis of the environment parameters, the resulting video plans can differ substantially, covering a range of possible outcomes.
@@
where $K$ denotes the distance function. The rejection module then selects the plan with the largest distance to past failures,
@@ 4.3 Action Module
The final step is to convert a selected video plan into executable low-level actions for the robot. We follow the idea of dense object- or wrist-tracking from AVDC (Ko et al., 2024): the predicted video frames are used to track either the object of interest or, in tasks such as bar picking, the robot’s wrist when object tracking is insufficient. The resulting trajectory is then translated into control commands through simple heuristic policies such as grasp strategies or trajectory-following controllers. This training-free approach enables direct execution of video plans without requiring an additional learned policy.
@@ 5 Experiment
5.1 Meta-World System Identification Suite
Figure 4: Qualitative results of adapted video plans. We show qualitative examples of adaptation trials for each task in the Meta-World Identification Suite to illustrate the effectiveness of the proposed method. The framework generates new plans based on previously failed interaction and video plans.
Existing manipulation benchmarks primarily emphasize single-episode task completion while assuming full observability, overlooking the rich information available through interaction. Consequently, they fail to capture challenges posed by environmental variation, such as determining whether a door should be pushed or pulled, that humans frequently encounter in the real world and can solve with ease. Therefore, in this work, we propose the Meta-World System Identification Suite to evaluate online adaptation under unknown system parameters $\theta$. It combines an offline experience dataset with environments exhibiting hidden parameters that agents must infer through interaction. All environments act in the same 4-dimensional positional control action space (dx,dy,dz,g)(dx,\,dy,\,dz,\,g), where the first three components are end-effector displacements and $g$ controls the gripper. Figure 4 shows the set of 5 tasks, including object manipulation (e.g., push bar, pick bar) and puzzle solving (e.g., turn faucet), test adaptability to randomized centers of mass, uncertain interaction modes, and external forces. We provide more details for these tasks in Appendix B.
@@ 5.1 Meta-World System Identification Suite
Turn faucet. In this task, the robot must rotate a faucet. The faucet can be rotated either clockwise or counterclockwise.
@@ 5.2 Evaluation Metrics
Second, instead of only evaluating task success, we also assess the effectiveness of the replanning mechanism by comparing the replanned videos to ground-truth interaction videos. To quantify this similarity, we use standard video metrics: PSNR, LPIPS (Zhang et al., 2018), CLIP-score, and DINO-score.
@@
5.3 Implementation and Baselines
We evaluate our framework on the 5 tasks in the Meta-World System Identification Suite. For our model, the Video Plan Generator is trained on the videos in the offline experience dataset $\mathcal{D}$. Each task is evaluated over 400 trials. We compare our approach with the original AVDC implementation for video planning, which does not online adapt, variants of our method that use different context encoders $E$, and other reinforcement learning baselines.
AVDC (Ko et al., 2024). The baseline with only the Action Module, i.e., , without the Retrieval Module and Rejection Module introduced in this paper.
Ours. Our proposed method with $E$ based on DINOv2 feature as described in Section 4.1.
Ours (Refine). A variant of our method that refines state embeddings by sampling from the multivariate Gaussian distribution estimated from $\mathcal{D}$, as described in Section 4.1.
BC. A state embedding-conditioned behavior cloning baseline where the policy takes in state and state embedding directly. The state embedding is obtained with $E$ based on DINOv2.
CQL. (Kumar et al., 2020) A state embedding-conditioned actor-critic framework. The Q-function takes in state, action, and state embedding, while the Policy takes in state and state embedding, which is obtained with $E$ based on DINOv2.
@@ 5.4 Results
Our variants. Among the variations of our method, using the Rejection Module (Ours) alone to generate state embeddings achieves the best overall performance. In contrast, the refinement process introduced in Section 4.1 alone, Ours (Refine), yields similar or slightly worse performance but does not assume data availability at test time, i.e., do not require access to the dataset used to train the diffusion model. Additionally, its formulation as a learned model offers the potential for improved generalization.
@@
Method PSNR $\uparrow$ LPIPS $\downarrow$ CLIP $\uparrow$ DINO $\uparrow$
AVDC 19.426 0.271 0.870 0.705
Ours (R3M) 19.632 0.261 0.875 0.705
Ours (CLIP) 19.584 0.262 0.871 0.707
Ours (DINOv2) 19.817 0.255 0.882 0.717
Video plan evaluation. In Table 2, we report the evaluations of video plans generated by the baseline (AVDC) and our methods with various context encoders $E$, including R3M (Nair et al., 2022), CLIP (Radford et al., 2021), and DINOv2 (Oquab et al., 2024). Our method produces significantly higher-quality videos compared to AVDC. Our method with DINOv2 outperforms its counterpart using R3M or CLIP, validating our design choices.
(a) Rejection and retrieval modules. Both the Rejection and Retrieval Modules improve performance, and combining them yields the best results.
(b) Rejection module distance. We compare raw-pixel distance and DINOv2 embedding distance for rejection, where the former yields better performance.
(c) Number of generated video plans. We vary the number of video plans $n\in[1,5]$ and set $n=2$ to balance computational cost and performance gain.
Figure 5: Ablation studies. We evaluate variations of our method to assess the effectiveness of each component. The number of replans is normalized and aggregated across five tasks.
@@ 5.5 Ablation Studies
To investigate the effectiveness of the Rejection and Retrieval Modules, we compare our full method against variants that remove either the Rejection or the Retrieval component. Notably, the Ours w/o Retrieval baseline is conceptually related to recent work (Li et al., 2025), which samples candidate plans using a generator and selects among them based on feedback from previous interaction attempts. The result in Figure 5(a) shows the normalized number of replans aggregated across all tasks. Removing either component consistently degrades performance, highlighting the importance of the Rejection and Retrieval modules. The full method achieves the best overall performance, validating the effectiveness of our design.
@@
We compare several distance functions for the Rejection Module in Figure 5(b) to assess their impact on overall performance. Both the DINOv2-based feature distance and the raw-pixel distance provide effective discrepancy signals for rejection, with the raw-pixel distance performing slightly better in our experiments. While the Rejection Module plays a central role in incorporating discrepancy signals during planning, our framework does not rely on a specific distance function, and this choice can be flexibly substituted. Accordingly, we adopt the raw-pixel distance as the default baseline in the main results. Additional analysis is provided in Section D.7.
@@
We study the effect of the number of generated video plans used by the Rejection Module by varying this number from 1 to 5. Figure 5(c) reports performance aggregated across all tasks. Increasing the number of candidates from $n=1$ to $n=2$ (our default) reduces the number of replans by approximately $10\%$. Beyond this point, increasing the number of candidates further provides diminishing returns: the performance for $n\in\{2,3,4,5\}$ is nearly identical. To balance computational cost and performance gain, we choose $n=2$ for our method.
Figure 6: Real-world setup. A WidowX arm opening a door that can only be pulled or pushed open.
5.6 Evaluation on Real-world Data
@@
We collected 20 real-world videos (10 successes, 10 failures) to train the video model. We address the challenge of training on such a small dataset by adopting a joint training with simulation data: 40% of samples came from real-world data and 60% from the Meta-World System Identification Suite. To reduce overfitting to the first frames, which often leads to monotonic plans, we injected Gaussian noise into the first frames during inference. Furthermore, we set $n=4$ to promote diverse plan generation.
Figure 7: Real-world qualitative results. Given the first frame (a), the robot first attempts to push the door (b). Then, it gets stuck and fails to open the door (c). Finally, it generates a plan that pushes the door instead (d) by leveraging information from previous failures.
For simplicity, we evaluate the single-episode replanning success rate, conditioned on an initial failure—different from the main evaluation metric. A perfect random policy, which assumes knowledge of the correct modes (e.g., push or pull) but does not use past interaction, would succeed with 50% probability. Our method achieved 15/20 successes, compared to 8/20 for a naive video planner without using past information, demonstrating the effectiveness of our method. Figure 7 shows the qualitative result, and more details are described in Section C.5.
@@ 6 Conclusion
We presented a video planning framework that adapts to interaction-time uncertainties by implicitly encoding past experiences into the planning process. Our method filters failed plans, updates models online, and achieves robust adaptation without explicit system identification. Experiments on the Meta-World System Identification Suite confirm improved robustness over video-based and RL baselines, highlighting the potential of video as a foundation for adaptive decision-making in uncertain environments.
Impact Statements
This paper presents work whose goal is to advance the field of Robot Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
This work was supported by the Industrial Technology Research Institute, Taiwan, the National Science and Technology Council, Taiwan (114-2628-E-002 -021-), and the Taiwan Centers of Excellence. Shao-Hua Sun was supported by the Yushan Fellow Program by the Ministry of Education, Taiwan.
@@ A.1 Website and Code
We present extensive qualitative results and release our code in our website. Our codebase can be found here.
@@ Appendix B Benchmark Details of Simulator Tasks
The environments of the five tasks are shown in Figure 8.
B.1 Push Bar
The objective of this task is to push a bar towards the target. Physically, the bar is combined with a mass point and a mass-free bar. Therefore, if the robot does not push the bar exactly at the center of mass, torque is imposed on the bar, resulting in the rotation of the bar.
The bar has a length of 0.4 meters. For ease of data collection, we discretize the center of mass position during pretraining. The data distribution of the center of mass on the bar is shown in Figure 9. During evaluation, the C.M is randomly sampled from [0.18,0.18][-0.18,0.18].
0.18-0.180.180.18
Figure 9: Center of mass positions. We visualize the distribution of the center of mass for our bar tasks. Each point corresponds to a possible position of the center of mass.
The visualization of this task is shown in Figure 15.
B.2 Pick Bar
The objective of this task is to grasp a bar towards the target. Physically, like the push-bar mentioned above, the bar is combined with a mass point and a mass-free bar. Therefore, if the robot does not grasp the bar exactly at the center of the mass, torque is imposed on the bar, resulting in the tilt of the bar.
The properties of the bar in this task are the same as the bar in push-bar. And the distances between the center of mass and the geometric center of the bar in the pretraining stage are also the same as above.
The visualization of this task is shown in Figure 15.
B.3 Slide Brick
Note that the wooden slope has a relatively low friction coefficient and that the red slope has a relatively high friction coefficient. Therefore, the brick can slide down and accelerate in the wooden slope and decelerate on the red slope. The friction coefficient of the wooden slope is fixed and is set to 0, while the friction coefficient of the red slope varies. During training, we discretize the friction coefficient as $\mu\in\{0.24,0.25,0.26,\dots,0.40\}$ in our dataset $\mathcal{D}$. During evaluation, the friction coefficient is sampled from $[0.24,0.40]$.
The visualization of this task is shown in Figure 15.
B.4 Open Box
The visualization of the task is shown in Figure 15.
B.5 Turn Faucet
The visualization of the task is shown in Figure 15.
@@ C.3 Action module
C.4 Data collection process
In simulation, we use scripted policies to collect offline interaction data. Specifically, the policy takes as input both the current observation and a “guessed” system parameter. For example, in the Bar Picking task, the guess is a one-dimensional scalar representing the position of the center of mass relative to the bar. If this guess matches the true underlying parameter of the environment (e.g., , both are centered), the interaction results in a successful video; otherwise, it produces a failed interaction. These outcomes correspond directly to the success signal $s_{i}$ described in Section 3.
For each object, we perform multiple interactions with different guesses, generating several failed trials, followed by an interaction using the correct guess, which produces a successful trial. We then assign all trials from the same object the same object ID $o_{i}$ and terminate the interaction loop for this object. The policy then starts interacting with a new object by randomizing the initial configuration and hidden parameters. As for the real-world experiment, we simply used tele-operation to replace the guess-augmented scripted policy.
C.5 Real-world experiment
Figure 7 shows the qualitative rollout for our real-world experiment. We followed most of the hyperparameters as used in the Meta-World System Identification Suite. Additionally, we changed $n=4$ and injected Gaussian noise of standard deviation 0.35 into the first frames with pixel values normalized to $[0,1]$ to mitigate the first-frame bias issue.
@@ D.1 State Embedding Feature Space
Figure 16 visualizes the feature space introduced in Section 4.1, using DINOv2 (Oquab et al., 2024) as the encoder. The plots are generated by applying t-SNE to sets of state embeddings from each environment, where for continuous modes we uniformly sample four values for visualization. Our method yields a well-structured feature space in which different modes can be separated without additional data-processing, facilitating more effective downstream learning.
@@
D.2 Runtime analysis
Table 4: Computation time (in seconds) averaged over 100 trials.
Component AVDC Ours w/o. Refine Ours
Retrieve N/A 0.487 $\pm$ 0.008 11.876 $\pm$ 0.036
DDIM 0.777 $\pm$ 0.001 1.573 $\pm$ 0.004 1.599 $\pm$ 0.006
Reject N/A 1.276 $\pm$ 0.075 1.230 $\pm$ 0.068
Optical Flow 1.564 $\pm$ 0.021 1.661 $\pm$ 0.030 1.761 $\pm$ 0.028
Total 2.341 $\pm$ 0.021 4.997 $\pm$ 0.081 16.466 $\pm$ 0.082
As shown in Table 4, we measured the average computation time over 100 iterations on a machine with an Intel Xeon W-2255 CPU and an NVIDIA RTX 3080 Ti GPU. Each iteration includes embedding retrieval, video generation (DDIM), plan selection (reject), and action decoding via optical flow. Although our full pipeline is slower per replan, this does not imply worse time-to-success. A more relevant metric for practical deployment is wall-clock time until the first successful execution, rather than latency per planning step. For example, in the Push Bar task, AVDC requires an average of 6.83 replans before success, with approximately 2.3 seconds per replan. Our full method requires fewer replans (4.26 on average), but each replan is slower (approximately 16.5 seconds).
In addition, when a plan fails, execution requires approximately 25 seconds to terminate before replanning. Taking this into account, the expected time to first success is:
  • AVDC: 6.83 × (2.3 + 25) = 186.5 seconds
  • Ours: 4.26 × (16.5 + 25) = 176.8 seconds
Despite a higher per-replan cost, our method reaches the first success faster overall due to the substantially lower failure rate.
Moreover, our computational cost is amortizable, while execution time is not. With better hardware or a more accelerated algorithm, the refinement overhead can be reduced, whereas the real-world cost of a failed execution remains fixed. This means the time advantage of our method naturally increases with improved computing infrastructure.
Finally, our method provides an additional benefit that AVDC does not: once a state embedding is identified for a particular object, it can be reused in future interactions. This allows the agent to succeed with higher probability on subsequent trials without repeating the identification process, whereas AVDC would again incur the full failure cost of approximately 186.5 seconds to reach its next success.
D.3 Training Data and Scaling Effects.